# RHuman: Real-Time 3D Human Appearance Rendering from a Single Image

### 10 Dec 2023 v1

@@ 6 Results on 1K Resolution
Our method is trained and tested on a resolution of . We also explore the performance of the method at 1K resolution. Specifically, we have remade a 1K dataset and trained our model on it. Fig. 10 shows the results of our method at high resolution. It can be seen that high-resolution images effectively improve the clarity of the results and enhance the visual effect of the results.
6.1 Application
We apply our method to real-time monocular human novel view synthesis, operates as follows: Images are captured from the video stream using OpenCV [3] and preprocessed to a resolution. We then apply RVM [11] for human mask extraction, effectively removing the background, and use HMR2.0 [8] for SMPL model estimation. Subsequently, the FOF-SMPL [4] is utilized to infer the occupation field. CUDA is employed to accelerate the rendering of depth, normal, and IUV maps. The final high-fidelity novel view images are rendered in real-time using the rotation matrix and the previously processed data. Our pipeline, implemented using TensorRT on a single RTX-3090 GPU, achieves a performance of 24+ FPS, with potential for further enhancement via additional GPUs. Fig. 11 and Fig. 12 showcase some of the real-time reconstruction results.
| & SMPL | information | ||||
| Time(ms) | 4.21 | 10.59 | 8.77 | 0.93 | 13.13 |
Tab. 5 details the execution time for each component of our pipeline. Currently, we implement our pipeline serially, achieving an overall runtime of approximately 37ms. We have combined the neural network and tensor multiplication into a single component, optimized using TensorRT for enhanced acceleration. Moreover, there is potential for substantially increasing the pipeline’s processing speed through parallel implementation of each component.
@@ 7 Conclusion and Discussion
Conclusion. We propose a novel method, combining the strengths of implicit texture field and explicit neural rendering, for real-time inference and rendering of realistic 3D human appearance from a monocular RGB image. Our approach introduces a depth mapping between the novel and source views, which reduces depth ambiguities in rendering and enables high-fidelity color reconstruction. The experimental results show that the proposed method achieves state-of-the-art performance on both synthetic data and real-world environment.
Limitation. Although our method makes full use of the 3D geometry information in the reconstructed data, there may still be artifacts caused by inaccurate geometry estimation when the angle changes are too large. In addition, while our method can provide relatively consistent results when synthesizing novel views with changing view directions, temporal inconsistencies may still happen because the feature warp varies slightly from frame to frame.

### 21 Mar 2024 v2

@@ 4.4 Results on 1K Resolution
Our method in the paper was trained and tested on a resolution of . We also explored the performance of the method at a higher resolution of 1K 1K. Specifically, we created a 1K dataset and trained our model on it. Fig. 12 presents the results of our method at this higher resolution. It can be seen that the use of high-resolution images significantly improves the sharpness of the results and enhances visual effects.
4.5 Failure Cases
Although our method makes full use of the 3D geometry information in the reconstructed data, there may still be artifacts caused by inaccurate geometry estimation. In addition, due to the absence of an explicit segmentation of skin parts and clothing parts, there are cases where the network may produce inaccurate color estimations on specific body parts. Fig. 13 illustrates some examples of these failure cases. In the future, we plan to explore additional supervision methods to enhance the network’s ability to recognize human body parts more effectively and thereby improve the quality of the output images.
@@ 5 Conclusion and Discussion
Conclusion. Rendering 3D human appearance from a single image in real time is important for achieving holographic communication and immersive VR/AR. We propose a novel method, combining the strengths of implicit texture field and explicit neural rendering, for real-time inference and rendering of realistic 3D human appearance from a monocular RGB image. We propose a new representation, Z-map, to alleviate depth ambiguities in rendering and enable high-fidelity color reconstruction. We also design a consistency loss and a spatio-temporal fusion strategy to ensure the multi-view coherence and reduce the jittering phenomenon. Experimental results show that the proposed method achieves state-of-the-art performance on both synthetic data and real-world images.